# Unifying Lane-Level Traffic Prediction from a Graph Structural Perspective: Benchmark and Baseline

### v1

@@ Abstract
Traffic prediction has long been a focal and pivotal area in research, witnessing both significant strides from city-level to road-level predictions in recent years. With the advancement of Vehicle-to-Everything (V2X) technologies, autonomous driving, and large-scale models in the traffic domain, lane-level traffic prediction has emerged as an indispensable direction. However, further progress in this field is hindered by the absence of comprehensive and unified evaluation standards, coupled with limited public availability of data and code. This paper extensively analyzes and categorizes existing research in lane-level traffic prediction, establishes a unified spatial topology structure and prediction tasks, and introduces a simple baseline model, GraphMLP, based on graph structure and MLP networks. We have replicated codes not publicly available in existing studies and, based on this, thoroughly and fairly assessed various models in terms of effectiveness, efficiency, and applicability, providing insights for practical applications. Additionally, we have released three new datasets and corresponding codes to accelerate progress in this field, all of which can be found on https://github.com/ShuhaoLii/TITS24LaneLevel-Traffic-Benchmark.
@@ I Introduction
Traffic plays a crucial role in modern society, impacting urban planning, traffic management, environmental protection, and daily life. Consequently, traffic forecasting has emerged as a hot and key research area. Current studies span a range of scales from macro to micro, including city-level, regional-level, road-level, and lane-level predictions.
City-level predictions focus on utilizing transfer learning[1, 2] or multi-task learning[3] to solve urban-scale traffic issues such as congestion and accident hotspot analysis[4]. Regional-level traffic prediction employs various methods to enhance understanding of traffic patterns within specific areas[5, 6]. This involves analyzing traffic flow, travel times, and vehicle types in particular zones like commercial, residential, or industrial areas[7]. The most extensive research is in road-level traffic prediction, involving modeling the spatio-temporal dependencies of roads using a combination of GCN-based (Graph Convolutional Networks)[8, 9, 10], RNN-based (Recurrent Neural Networks)[11], and Attention-based mechanisms[12, 13]. These methods capture the dynamic temporal changes and spatial correlations in traffic flow, thereby improving the accuracy of traffic forecasts. In lane-level traffic prediction, the focus shifts to more granular traffic flow analysis. This includes real-time monitoring and forecasting of parameters like traffic volume and speed for individual lanes. The challenge at the lane level lies in handling higher frequency and finer-grained data while considering interactions and lane-changing behaviors among vehicles. This level of prediction is vital for the design and implementation of intelligent traffic systems, especially against the backdrop of rapidly advancing autonomous driving technologies and Vehicle-to-Everything (V2X) communication[14, 15]. Lane-level predictions not only enhance traffic efficiency but also reduce accidents and environmental impacts.
@@
Figure 1: The current state of public accessibility for codes and datasets in existing lane-level research studies. Note: ’Available’ here refers to the explicit mention of accessible code or dataset URLs in their respective research papers.
Although lane-level traffic prediction has undergone a lengthy research journey, its progress has been somewhat slower compared to more macroscopic levels of prediction, such as regional-level or road-level prediction.
The achievements in this field are relatively scattered and follow a development trend similar to road-level traffic prediction. Early research mainly relied on basic statistical methods and simple flow models to predict lane-level traffic[16, 17]. However, with technological advancements, the focus shifted towards applying more complex machine learning and deep learning methods[18, 19]. Despite this progress, the application of these advanced methods still faces numerous challenges, such as complex intersections on urban expressways and inconsistent numbers of lanes. Moreover, lane-level traffic prediction involves processing large amounts of high-frequency, high-dimensional data, increasing computational costs and imposing greater demands on data collection and processing.
Additionally, there are many unresolved issues in the field. For instance, how to rapidly and accurately respond to changes in lane traffic flow, and how to effectively integrate data from diverse sources and types to enhance prediction accuracy. Furthermore, the lane-level traffic prediction domain lacks a comprehensive overview or established benchmark standards, making it difficult for newcomers and practitioners to quickly grasp the core issues and latest developments in the field.
Moreover, as summarized in Figure 1, the majority of studies are conducted on private datasets, with neither their code nor datasets being made public [20, 21, 22, 23, 24, 19, 25]. This lack of sharing and openness severely limits the comparison, validation, and further improvement of research methods. Therefore, to drive progress in this field, there is an urgent need to establish public standard datasets and benchmark tests to encourage the research community to share codes and findings. This will not only improve the quality of research but also accelerate the advancement of the entire field.
@@
We commence with a detailed review and classification of current research in lane-level traffic prediction. This review emphasizes the development of spatial topology and the techniques for modeling spatial-temporal dependencies. Additionally, we conduct an in-depth analysis of the advantages and limitations of these methods. Building upon this groundwork, our approach to traffic prediction incorporates a macro perspective and the application of graph structures. We effectively establish lane-level network topologies using graph structure. Utilizing this framework, we develop an advanced, efficient solution and a novel baseline prototype model. We then extract data from three real-world lane networks, comprising two with regular lane configurations and one with irregular lane configurations. To enhance the evaluation of model performance, we introduce training cost as a new metric. This metric seeks a balance between model effectiveness and efficiency, both of which are vital in lane-level traffic prediction. It evaluates the models based on both prediction accuracy and training duration, ensuring a thorough assessment. Finally, we conduct extensive and impartial testing of existing lane-level traffic prediction models and our proposed methods using these datasets and the training cost metric. These tests confirm our model’s efficacy and highlight its potential in practical scenarios, particularly in managing complex traffic patterns and enabling efficient real-time predictions. Our research contributes novel insights and tools to the field of lane-level traffic prediction, establishing a robust foundation and direction for future scholarly and practical endeavors.
In summary, the primary contributions of this study are as follows:
  • Comprehensive Literature Review: We present the first extensive review and categorization of existing lane-level traffic prediction research, highlighting their strengths and weaknesses.
@@
  • Unified Topological Modeling and Baseline: We propose a unified topological modeling approach and predictive tasks for various lane-level traffic predictions, along with a simple baseline model tailored to these challenges.
  • Creation of Diverse Datasets: We extract and make public three unique datasets from real-world road networks, encompassing both regular and irregular lane configurations.
  • Improved Code Replication and Fair Benchmarking: We replicate codes that were previously unavailable in existing research and conduct thorough, equitable benchmarking of current and new models, emphasizing both their effectiveness and efficiency.
@@
The rest of this paper is organized as follows: In Section 2, we provide a comprehensive classification and analysis of existing work based on spatial topology construction and spatio-temporal dependency modeling. Subsequently, Section 3 defines a unified problem of lane-level traffic prediction. Motivated by these challenges, details of our proposed unified graph-structured approach and a simple baseline are introduced in Section 4. Next, in Section 5, we evaluate our model on three datasets using multiple metrics, presenting a benchmark for lane-level traffic prediction, including studies on effectiveness and efficiency, as well as long-term predictive performance. Finally, we conclude the paper and discuss the future implications of lane-level traffic prediction in Section 6.
@@
Figure 2: Categorization of existing lane level traffic prediction research based on spatial topology construction, spatial dependency modeling techniques, and temporal dependency modeling techniques.
II Literature Review
@@
Lane-level traffic prediction is an integral part of intelligent transportation systems. The challenge in this field lies in accurately capturing and understanding the complex spatio-temporal dependencies in traffic flow. Spatial dependencies refer to how traffic flow on different roads or lanes affects each other, while temporal dependencies focus on the patterns of traffic flow over time. These models usually need to predict traffic volume accurately in highly dynamic and multi-variable urban traffic networks, requiring not only the capture of complex traffic dynamics but also a rapid response to real-time changes in traffic conditions. As illustrated in Figure 2, based on existing research, we categorize the modeling approaches for spatial topology (modeling objectives), spatial dependency modeling techniques, and temporal dependency modeling techniques. By analyzing the core principles, strengths, and limitations of various modeling techniques, we aim to provide a comprehensive perspective for understanding the current state and future directions of lane-level traffic prediction.
@@ II-A Spatial Topology Construction
The construction of spatial topology is crucial for modeling objectives and significantly impacts spatio-temporal dependency modeling and prediction accuracy. Lane-level traffic prediction, as compared to road-level or city-level predictions, presents more complexity and variety in its modeling subjects. To better comprehend the dynamics of traffic at this level, as illustrated in Figure 3, lane-level traffic prediction can be categorized into four main types of modeling objects: Entire Lanes, Grid-Based Models, Intersection Approaches, and Lane Segments. A brief introduction to these categories is as follows:
@@ II-A1 Entire Lanes
Some research considers an entire lane as the modeling subject, focusing on the overall traffic flow of a single lane[20, 25]. The number of modeling subjects in this method usually equals the number of lanes. However, this approach may not effectively capture the interactions between lanes and the traffic flow patterns of the overall road network.
II-A2 Grid-based
Numerous studies conceptualize lane segments as grid-like structures, akin to pixels in an image, and employ Convolutional Neural Networks (CNN) for modeling[22, 26, 27].These studies, by taking into account the influence of adjacent lanes, typically achieve favorable outcomes. However, a significant limitation of this method is its inability to effectively model road segments that feature entrances and exits or vary in the number of lanes.
@@ II-A3 Intersection Lanes
Some studies focus on lane-level prediction for lanes at intersections[28, 23, 24, 18]. Due to their importance and the complexity of traffic flow, intersections are key areas for lane-level prediction. Modeling traffic flow at intersections can help improve traffic congestion management, optimize traffic signal control, and enhance driving safety. However, these studies often use road-level datasets and may lack detailed analysis of lane-level traffic flow, limiting their applicability in universal lane-level traffic prediction.
@@ II-A4 Lane Segments
This topological structure initially segments the road network into road segments and then further divides these into individual lane segments[21, 29, 19], It represents a more flexible topological structure, where even the lanes at intersection lanes can be considered as lane segments with distinct geographical positioning. This approach addresses the limitations inherent in grid-based methods, enabling a more comprehensive capture of the dynamic characteristics of lane-level traffic flow. However, this also implies an increase in the complexity of the modeling objects, posing greater challenges for spatial dependency modeling techniques.
@@ II-B Spatial Dependency Modeling
Spatial dependency modeling is critical in lane-level traffic prediction as it involves accurately capturing the interactions of traffic flow among different lanes and road segments. Overall, spatial dependency modeling can be categorized into several key technical approaches, including methods without specific spatial modeling[28], signal processing techniques[20, 21], and data-driven deep learning models[22, 26, 27, 23, 24, 29, 18, 19, 25]. Each category aims to tackle the issue of spatial dependency from various angles and methodologies, striving to enhance the accuracy of traffic flow predictions.
II-B1 Signal Processing Techniques
In the realm of signal processing techniques, the CEEMDAN-XGBoost model[20] employs the Complete Ensemble Empirical Mode Decomposition with Adaptive Noise(CEEMDAN) method to decompose traffic flow data, extracting spatial features, while the Fusion Deep Learning (FDL) model[21] uses entropy-based grey relational analysis to identify lane segments strongly correlated with the target lane’s traffic conditions. Both methods aim to enhance prediction accuracy by analyzing spatial correlations within the traffic network. However, their limitation lies in potentially not fully leveraging the global topological structure of the traffic network, which may lead to inadequacies in capturing more complex spatial dependencies.
In the field of deep learning, the approaches primarily fall into categories such as convolutional and graph convolutional methods, hybrid deep learning methods, attention-based methods, and multi-graph fusion and visible graph methods.
II-B2 Convolutional and Graph Convolutional Methods
These approaches use a combination of CNNs (Convolutional Neural Networks) and GCNs (Graph Convolutional Networks) to enhance the capture of spatial features. Models like TM-CNN[22] and GCN models based on data-driven adjacency matrices[23, 24, 29] are typical examples in this category. They effectively leverage the local feature extraction capability of CNNs and the global spatial relationship modeling ability of GCNs, thus efficiently capturing the complex spatial patterns of lane-level traffic flows. However, the primary drawback of convolution-based methods is their treatment of lane networks as image-like structures, which fails to model roads with varying numbers of lanes. Graph convolutional approaches are comparatively more flexible.
II-B3 Hybrid Deep Learning Methods
Models like MDL[26] combine different types of neural networks, such as ConvLSTM[27], convolutional layers, and dense layers, to capture the spatio-temporal features of traffic data. This hybrid approach offers a more comprehensive perspective in understanding the dynamics of lane-level traffic flow, especially in capturing temporal and spatial correlations. While these methods consider both temporal and spatial features, their typically complex structures may lead to higher computational costs and longer training times.
II-B4 Attention-based Methods
Include ST-AFN[18]and encoder-decoder-based spatio-temporal attention models[19], as well as attention mechanisms between multi-view graph structures[25]. These models, by introducing attention mechanisms, flexibly capture the dynamic relationships between lanes and temporal changes. The attention mechanism helps focus the model more on the spatial areas and time points most crucial for prediction. Despite the flexibility and focused capability provided by attention mechanisms, they may increase the complexity of the model, making training and fine-tuning more challenging.
In summary, while each of these deep learning methodologies offers unique advantages in modeling spatial dependencies in lane-level traffic prediction, they also present their own set of challenges, including complexity, computational demand, and the need for specialized knowledge in model development and tuning.

### v2

@@ Abstract
Traffic prediction has long been a focal and pivotal area in research, witnessing both significant strides from city-level to road-level predictions in recent years. With the advancement of Vehicle-to-Everything (V2X) technologies, autonomous driving, and large-scale models in the traffic domain, lane-level traffic prediction has emerged as an indispensable direction. However, further progress in this field is hindered by the absence of comprehensive and unified evaluation standards, coupled with limited public availability of data and code. In this paper, we present the first systematic classification framework for lane-level traffic prediction, offering a structured taxonomy and analysis of existing methods. We construct three representative datasets from two real-world road networks, covering both regular and irregular lane configurations, and make them publicly available to support future research. We further establishes a unified spatial topology structure and prediction task formulation, and proposes a simple yet effective baseline model, GraphMLP, based on graph structure and MLP networks. This unified framework enables consistent evaluation across datasets and modeling paradigms. We also reproduce previously unavailable code from existing studies and conduct extensive experiments to assess a range of models in terms of accuracy, efficiency, and applicability, providing the first benchmark that jointly considers predictive performance and training cost for lane-level traffic scenarios. All datasets and code are released at https://github.com/ShuhaoLii/LaneLevel-Traffic-Benchmark.
@@ I Introduction
Lane-level traffic prediction has gradually become an essential component of multi-granularity traffic prediction with the rapid development of intelligent vehicles and refined urban management. In the hierarchical structure of traffic prediction, city-level studies primarily address large-scale urban transportation issues using methods such as transfer learning [1, 2] and multi-task learning [3], providing foundational support for urban planning and smart city infrastructure [4]. Regional-level prediction focuses on enhancing the understanding of traffic patterns in specific areas [5, 6], such as analyzing traffic flow, travel time, and vehicle types in commercial, residential, or industrial zones [7]. Road-level traffic prediction is one of the key focuses of current research. Its core lies in leveraging models such as graph convolutional networks (GCNs) [8, 9, 10], recurrent neural networks (RNNs) [11], and attention mechanisms to capture the spatio-temporal dependencies at the road level [12, 13]. This line of research aims to accurately predict traffic states, alleviate congestion, and identify high-risk accident zones.
As illustrated in Figure 1, lane-level traffic prediction further refines the research perspective by concentrating on real-time monitoring and prediction of parameters at the single-lane level, such as traffic flow and vehicle speed. Compared to other levels, lane-level prediction requires processing higher-frequency and more granular data while simultaneously considering vehicle interactions and lane-changing behaviors. Research at this level is particularly significant for optimizing intelligent transportation systems [14, 15]. For example, it can assist autonomous vehicles in selecting lanes with lower traffic volumes, reduce the need for frequent lane changes by freight vehicles, and improve traffic safety. Additionally, it provides critical technical support for the dynamic adjustment of traffic signals and flexible management of reversible lanes.
Figure 1: Examples of comparisons between regional/road-level and lane-level traffic prediction and their supported application scenarios: lane-level traffic prediction plays a critical role in more refined traffic management and applications.
Despite having a relatively long research history, lane-level traffic prediction remains underdeveloped compared to other granularities of traffic prediction. Existing research findings are relatively fragmented, resembling the early development of road-level traffic prediction. Initial studies primarily relied on basic statistical methods and simple flow models. With technological advancements, more complex machine learning, and deep learning methods have gradually been introduced into the research. However, the application of these methods still faces numerous challenges. For instance, issues such as inconsistent lane counts and complex multi-branch intersections on urban expressways significantly increase modeling difficulty.
Although lane-level traffic prediction has accumulated some research, its development has significantly lagged behind that of other granularities of traffic prediction. Existing findings are relatively scattered, following a research trajectory similar to the early stages of road-level traffic prediction. Early studies primarily relied on basic statistical methods and simple flow models [16, 17], which, while computationally inexpensive, struggled to effectively capture the complex interactions and dynamic characteristics between lanes. Moreover, traditional statistical models often assumed linear traffic flow relationships, overlooking the highly nonlinear and time-varying nature of real-world traffic flows, resulting in limited prediction accuracy. With advancements in technology, more complex machine learning and deep learning methods have been gradually introduced [18, 19]. These methods leverage the strong modeling capabilities of neural networks to capture the spatio-temporal dependencies in lane-level traffic flows. However, their practical application still faces numerous challenges. For instance, inconsistencies in lane counts and the complex multi-branch layouts of urban expressways significantly increase modeling difficulty. Additionally, lane-level traffic prediction requires handling massive, high-frequency, high-dimensional data, which not only significantly raises computational costs but also imposes higher demands on data collection and processing capabilities.
@@
TABLE I: The current state of public accessibility for codes and datasets in existing lane-level research studies.
Models Code Available Dataset Available
Cat-RF-LSTM [20]
CEEMDAN-XGBoost [21]
STMGG [22]
TM-CNN [23]
MDL [24]
CNN-LSTM [25]
HGCN [26]
DGCN [27]
ST-AFN[18]
FDL [28]
GCN-GRU [29]
STA-ED [19]
  • Note: ’Available’ here refers to the explicit mention of accessible code or dataset URLs in their respective research papers.
Several unresolved issues remain in this field, such as how to rapidly and accurately respond to changes in lane traffic conditions, and how to integrate lane traffic prediction across different scenarios to achieve a universal and unified solution for lane-level prediction. Furthermore, there is currently no comprehensive review or benchmark research, making it difficult for new researchers entering the field to quickly grasp its core problems and latest advancements. As shown in Table I, most existing studies are based on private datasets, with relevant code and datasets not publicly available [21, 28, 23, 26, 27, 19, 22]. This lack of sharing and openness severely limits the comparison, validation, and improvement of research methods. To advance the field of lane-level traffic prediction, it is imperative to establish public standardized datasets and benchmarks and to encourage the research community to share code and data resources openly. This would not only improve the quality of research but also accelerate technological progress across the entire field.
@@
We commence with a detailed review and classification of current research in lane-level traffic prediction. This review emphasizes the development of spatial topology and the techniques for modeling spatial-temporal dependencies. Additionally, we conduct an in-depth analysis of the advantages and limitations of these methods. Building upon this groundwork, our approach to traffic prediction incorporates a macro perspective and the application of graph structures. We effectively establish lane-level network topologies using graph structure. Utilizing this framework, we develop an advanced, efficient solution and a novel baseline model. We then extract data from three real-world lane networks, comprising two with regular lane configurations and one with irregular lane configurations. To enhance the evaluation of model performance, we introduce training cost as a new metric. This metric seeks a balance between model effectiveness and efficiency, both of which are vital in lane-level traffic prediction. It evaluates the models based on both prediction accuracy and training duration, ensuring a thorough assessment. Finally, we conduct extensive and impartial testing of existing lane-level traffic prediction models and our proposed methods using these datasets and the training cost metric. These tests confirm our model’s efficacy and highlight its potential in practical scenarios, particularly in managing complex traffic patterns and enabling efficient real-time predictions, as well as identifying critical lane segments, recurrent congestion points, and anomalous traffic patterns through fine-grained modeling. Our research contributes novel insights and tools to the field of lane-level traffic prediction, establishing a robust foundation and direction for future scholarly and practical endeavors.
In summary, our main contributions can be outlined as follows:
  • We introduce the first systematic classification framework for lane-level traffic prediction and provide a comprehensive analysis of the technical characteristics, advantages, and limitations of existing methods.
@@
  • We propose a unified topology modeling approach, clarify the tasks and challenges of lane-level traffic prediction, and design a simple yet effective baseline model to address these challenges.
  • We extract and construct three reliable datasets from two real-world road networks, including datasets with regular and irregular lane counts, to enhance the diversity of testing scenarios. These datasets are made publicly available to the research community.
  • We reproduce and release previously unavailable code from existing studies, conduct extensive and fair testing of both existing models and the proposed baseline model, and evaluate their performance using dual metrics of effectiveness and efficiency.
@@
The rest of this paper is organized as follows: In Section II, we provide a comprehensive classification and analysis of existing work based on spatial topology construction and spatio-temporal dependency modeling. Subsequently, Section III defines a unified problem of lane-level traffic prediction. Motivated by these challenges, details of our proposed unified graph-structured approach and a simple baseline are introduced in Section IV. Next, in Section V, we evaluate our model on three datasets using multiple metrics, presenting a benchmark for lane-level traffic prediction, including studies on effectiveness and efficiency, as well as long-term predictive performance. Finally, we conclude the paper and discuss the future implications of lane-level traffic prediction in Section VI.
@@
Figure 2: Categorization of existing lane-level traffic prediction research based on spatial topology construction, spatial dependency modeling techniques, and temporal dependency modeling techniques.
II Literature Review and Classification
@@
Lane-level traffic prediction is an integral part of intelligent transportation systems. The challenge in this field lies in accurately capturing and understanding the complex spatio-temporal dependencies in traffic flow. Spatial dependencies refer to how traffic flow on different roads or lanes affects each other, while temporal dependencies focus on the patterns of traffic flow over time. These models usually need to predict traffic volume accurately in highly dynamic and multi-variable urban traffic networks, requiring not only the capture of complex traffic dynamics but also a rapid response to real-time changes in traffic conditions. As illustrated in Figure 2, based on existing research, we categorize the modeling approaches for spatial topology (modeling objectives), spatial dependency modeling techniques, and temporal dependency modeling techniques. Through Table II, we provide a detailed analysis of the core principles, advantages, and limitations of various modeling techniques, and we evaluate the strengths and weaknesses of existing research from three dimensions: flexibility, robustness, and efficiency. Specifically, flexibility assesses the applicability and transferability of the model, robustness evaluates the model’s ability to handle high dynamics and sudden changes, and efficiency focuses on the technical implementation and the scale of training parameters. Our goal is to offer a comprehensive perspective on the current state and future development directions of lane-level traffic prediction.
TABLE II: Overall comparison of the strengths and weaknesses of existing lane-level traffic prediction models across three dimensions.
Models Flexibility Robustness Efficiency
Cat-RF-LSTM [20]
CEEMDAN-XGBoost [21]
FDL [28]
STMGG [22]
TM-CNN [23]
MDL [24]
CNN-LSTM [25]
HGCN [26]
DGCN [27]
ST-AFN[18]
GCN-GRU [29]
STA-ED [19]
@@ II-A Spatial Topology Construction
The construction of spatial topology determines the modeling objectives and is a critical factor influencing the modeling of spatio-temporal dependencies and prediction accuracy. Compared to coarse-grained traffic prediction at the road or city level, lane-level traffic prediction requires capturing finer-grained dynamic data, significantly increasing the complexity and diversity of modeling targets. To systematically summarize existing research, we classified lane-level traffic prediction modeling objects into four main types based on different modeling requirements and application scenarios, as shown in Figure 3: Entire Lanes, Grid-based Models, Intersection Lanes, and Lane Segments. Below is a detailed explanation and analysis of the characteristics of each type:
@@ II-A1 Entire Lanes
The entire lane modeling approach treats each lane as a whole modeling object. Input data typically include overall lane-level traffic states, such as traffic flow, average speed, and lane density. The output targets are traffic predictions for each lane, such as future total traffic flow or average speed [21, 22]. A key advantage of this method is its simplicity: the number of modeling objects corresponds directly to the number of lanes, significantly reducing model complexity and improving computational efficiency. This approach is particularly suitable for scenarios with uniform traffic characteristics and minimal inter-lane interactions, such as unidirectional lanes on highways. However, it fails to capture interactions between lanes, making it less effective for complex segments like intersections or densely interactive urban roads. Consequently, its ability to model overall road network traffic patterns is limited.
II-A2 Grid-based Models
Grid-based methods divide lanes into fixed-sized uniform grids, treating each grid as a unit akin to an image pixel, allowing CNNs to extract local spatial features [23, 24, 25]. Input data typically includes traffic state information within grids, such as traffic density and local speed, while output targets are traffic predictions for each grid, such as future traffic flow or congestion levels. These methods perform well in simple road structures, effectively capturing dependencies between adjacent grid units. However, fixed grid divisions cannot adapt to roads with varying numbers of lanes, such as those with entrances or exits. This limitation reduces their effectiveness in dynamically changing traffic scenarios.
@@ II-A3 Intersection Lanes
Intersection modeling methods focus on the most complex areas of traffic flow and serve as a crucial approach for modeling key traffic nodes. In both road-level and lane-level traffic prediction, intersections remain one of the most challenging research subjects. Road-level intersection prediction typically relies on aggregated road-level traffic flow information, incorporating overall traffic movement patterns or leveraging traffic signal control to enhance the accuracy of regional traffic forecasts [30, 31]. In contrast, lane-level intersection modeling requires more granular input data, including lane-specific traffic volume, speed, vehicle turning behavior, and traffic signal control states. The output targets encompass lane-level traffic distribution prediction, traffic signal optimization parameter computation, and congestion evaluation [20, 26, 27, 18]. These methods play a significant role in optimizing traffic signal control, alleviating congestion, and enhancing driving safety. However, most current lane-level studies still predominantly depend on road-level datasets, failing to fully capture the fine-grained dynamic characteristics of intersections. The lack of high-resolution data and the limited adaptability of models in capturing cross-lane traffic flows constrain prediction accuracy, thereby affecting their applicability in dynamic and evolving traffic environments.
@@ II-A4 Lane Segments
Lane segment modeling divides each lane into multiple segments, treating each segment as an independent modeling unit, providing a more flexible spatial topology for complex traffic scenarios [28, 29, 19]. Input data typically include segment-level metrics such as flow, speed, density, and interaction features between adjacent segments. Output targets include segment-level flow predictions, congestion trends, or dynamic relationships between segments. This method is especially suited to multi-entry and multi-exit roads and other complex structures, capturing finer-grained dynamics. For example, in intersection traffic modeling, each lane segment can serve as an independent modeling object, significantly improving adaptability to dynamic scenarios. However, the increased number of modeling objects imposes higher computational demands and requirements for spatial dependency modeling techniques, potentially resulting in high computational costs in high-frequency, real-time traffic predictions.
II-A5 Comprehensive Analysis
Each existing spatial topology modeling method has its specific advantages and use cases, but also significant limitations. The entire lane modeling approach is simple and efficient, reducing model complexity and improving computational efficiency. However, it fails to effectively capture interactions between lanes, making it less suitable for complex traffic scenarios. Grid-based methods perform well in extracting local spatial features and are effective in simple road structures. Nevertheless, their reliance on fixed grid divisions limits their ability to adapt to dynamic topological changes, such as roads with varying lane counts. Intersection lane modeling provides a targeted approach for predicting traffic in critical areas with complex flows, such as intersections. While it enhances prediction accuracy and traffic signal optimization, its reliance on coarse data and limited ability to model fine-grained cross-lane interactions restricts its broader applicability. Lane segment modeling, on the other hand, excels at capturing detailed dynamics and adapting to complex road structures, making it particularly effective in scenarios with multi-entry and multi-exit roads. However, the significantly increased number of modeling objects leads to higher computational demands, limiting its feasibility for high-frequency, real-time applications.
These limitations underscore the need for a unified spatial topology structure that can accommodate diverse scenarios and flexibly adapt to dynamic changes. Such a structure should integrate various modeling objects, including entire lanes, intersection lanes, and lane segments, within a single framework to enhance the generalizability and practicality of lane-level traffic prediction models.
@@ II-B Spatial Dependency Modeling
Spatial dependency modeling is a crucial component of lane-level traffic prediction, aiming to accurately capture the complex interactions between lanes and road segments. Compared to traditional road-level prediction, lane-level traffic prediction deals with higher-resolution dynamic characteristics and diverse traffic scenarios. The design of efficient spatial dependency modeling techniques directly determines the performance of prediction models. Based on their technical principles, existing spatial dependency modeling methods can be broadly categorized into the following types: methods without specific spatial modeling [20], traditional methods based on signal processing [21, 28], and advanced methods based on deep learning [23, 24, 25, 26, 27, 29, 18, 19, 22].
II-B1 Methods Without Specific Spatial Modeling
These methods do not explicitly model the topology of the traffic network, treating lane-level flow prediction as a time series problem. Regression models (e.g., linear regression, support vector machines) or tree-based models (e.g., random forests, gradient boosting trees) are commonly employed to model lane flow directly based on historical data [20]. While computationally efficient, these methods rely solely on historical data from individual lanes and neglect spatial dependencies between lanes. Consequently, they perform poorly in capturing local interactions and global topological relationships. Their predictive performance is particularly limited in complex scenarios, such as multi-lane interactions or competitive flows at intersections.
II-B2 Signal Processing Techniques
Traditional methods based on signal processing use mathematical decomposition and statistical analysis to extract local spatial dependencies between lanes. For instance, the CEEMDAN-XGBoost model applies Complete Ensemble Empirical Mode Decomposition with Adaptive Noise (CEEMDAN) to decompose lane flow data into components of different frequencies for feature extraction and prediction [21]. Similarly, the FDL model utilizes entropy-based grey relational analysis to identify lane segments highly correlated with the target lane flow[28].
These methods excel at analyzing local lane associations and perform well in structured scenarios like unidirectional highways. However, they rely on static feature extraction and lack adaptability to dynamic topological changes. Moreover, their inability to capture global dependencies in complex road networks—such as those with multiple entrances or dynamically changing lane topologies—limits their applicability in more complex scenarios.
II-B3 Deep Learning-Based Methods
Deep learning models adopt data-driven approaches to learn complex nonlinear relationships in lane-level traffic flows, significantly enhancing predictive performance. These methods can be further divided into the following categories:
Convolutional Methods: CNNs and GCNs are two mainstream approaches for spatial modeling. CNNs excel at extracting local features from regular grid structures, while GCNs flexibly handle non-Euclidean structures like lane networks. For example, the TM-CNN model partitions lane flow data into uniform grids and uses CNNs to extract local dependencies [23], while GCN models leveraging data-driven adjacency matrices capture global spatial dependencies between lanes [26, 27, 29, 32, 33].
While CNN-based methods perform well in modeling interactions among adjacent lanes, they assume a regular grid structure, limiting their flexibility. On the other hand, GCNs can adapt to complex topologies but incur high computational costs during matrix operations. Their efficiency and adaptability in multi-entry or irregular topological scenarios still require improvement.
Hybrid Deep Learning Methods: Hybrid models integrate multiple network structures (e.g., convolutional layers, recurrent neural networks, dense layers) to capture both spatio-temporal dependencies. For instance, the MDL model combines ConvLSTM and dense layers to extract temporal dynamics and spatial dependencies through convolutional operations [24, 25]. These models are well-suited for modeling multi-lane interactions in complex road networks. However, their high computational complexity and resource demands make them less suitable for real-time applications with stringent latency requirements.
Attention-based Methods: Attention mechanisms dynamically assign weights to emphasize important spatio-temporal relationships[19, 22]. For example, the ST-AFN model combines spatial and temporal attention mechanisms to flexibly capture complex dynamic interactions between lanes [18]. These methods excel in highlighting critical interactions, such as lane-changing behaviors or entry flows. However, introducing attention mechanisms significantly increases model complexity, posing challenges for model training and parameter tuning.
II-B4 Comprehensive Analysis
The above classification reveals the strengths and limitations of various approaches to spatial dependency modeling. Methods without specific spatial modeling fail to capture topological relationships between lanes, limiting their ability to model complex interactions. Signal processing techniques are effective for local dependency modeling but struggle in dynamic scenarios. Deep learning-based models address these shortcomings with data-driven adaptability, achieving higher performance in complex lane networks. Within deep learning-based approaches, differences in effectiveness are also apparent. Convolution-based and graph convolution-based methods excel at capturing local and global dependencies but require improved adaptability for dynamic scenarios. Hybrid models offer comprehensive spatio-temporal feature extraction but are computationally expensive. Attention mechanism-based methods enhance the focus on critical dependencies but face limitations in scalability due to their complexity.
To address these challenges, future research should aim to balance modeling flexibility, computational efficiency, and real-time adaptability, ensuring robust performance across diverse traffic scenarios.